# MonoScale: Scaling Multi-Agent System with Monotonic Improvement

### v1

@@ Abstract
In recent years, LLM-based multi-agent systems (MAS) have advanced rapidly, using a router to decompose tasks and delegate subtasks to specialized agents. A natural way to expand capability is to scale up the agent pool by continually integrating new functional agents or tool interfaces, but naive expansion can trigger performance collapse when the router cold-starts on newly added, heterogeneous, and unreliable agents. We propose MonoScale, an expansion-aware update framework that proactively generates a small set of agent-conditioned familiarization tasks, harvests evidence from both successful and failed interactions, and distills it into auditable natural-language memory to guide future routing. We formalize sequential augmentation as a contextual bandit and perform trust-region memory updates, yielding a monotonic non-decreasing performance guarantee across onboarding rounds. Experiments on GAIA and Humanity’s Last Exam show stable gains as the agent pool grows, outperforming naive scale-up and strong-router fixed-pool baselines.
@@ 1 Introduction
Multi-agent systems (MAS) built on large language models (LLMs) are emerging as a powerful paradigm for solving complex tasks (Chen et al., 2025; Yang et al., 2024; Guo et al., 2024). Unlike a single model that produces an answer end-to-end, an MAS typically includes a router that decomposes an input problem, dispatches subtasks to specialized agents (Hu et al., 2025; Yang et al., 2025a; Yue et al., 2025)(e.g., retrieval, code execution, planning, tool use, or domain experts), and then aggregates their outputs into a final response. This modular “divide–and–conquer” structure enables flexible collaboration and makes MAS particularly effective for multi-step workflows, tool-augmented reasoning, and long-horizon task execution.
@@
Existing approaches for improving multi-agent coordination broadly fall into two lines, yet neither fully addresses the central challenge of open-ended MAS scaling. The first line focuses on a static agent pool, assuming that the agent set and its capability profile remain stable between training and deployment (Yue et al., 2025; Zhang et al., 2024; Liu et al., 2025b); the research emphasis is then on optimizing routing policies, improving division of labor, or planning more efficiently within a fixed pool. While effective in static settings, these methods typically lack a controlled update mechanism that can prevent performance regressions when new agents are continually added, and they do not provide a unified formalization for online updating under continual agent augmentation in open-ended systems. In parallel, some recent work has started to examine dynamic expansion in LLM routing (Wang et al., 2025; Zhang et al., 2025b), where the system routes across an enlarging set of available experts over time. However, scaling an MAS is substantially harder: newly integrated agents are far more heterogeneous (tools, memory, skills, roles, and interfaces), and their quality can be highly uneven in practice (brittle tool calls, API failures, retrieval noise, or unstable behaviors). As a result, the router faces a sharper cold start problem, and a fixed onboarding suite to “trial-run” the new agent is often insufficient: (i) it cannot be adapted to each agent’s unique boundaries and failure modes; and (ii) it may under-cover the true deployment distribution, creating a risk of agents that appear usable during onboarding but still collapse after real deployment. In short, what dynamic MAS scaling truly requires is an adaptive (agent-conditioned) and conservative update protocol, so that end-to-end performance can improve stably as the agent pool grows.
@@
To bridge this gap, we propose an update framework for sequentially expanding open-ended MAS. Whenever a new agent is introduced, rather than waiting for real user traffic to surface failures, we proactively customize a small set of familiarization tasks tailored to the new agent and use them to collect controlled feedback. Our mechanism has two stages: (i) active familiarization via task customization, where tasks are synthesized to probe the new agent’s likely strengths, interface constraints, and failure modes, enabling the router to learn when to use / when not to use the agent; and (ii) router update via memory, where we update the router’s policy by maintaining an auditable text memory distilled from experience. Concretely, the memory captures not only patterns behind successful orchestration, but also lessons from failed worker calls and mis-routings; we abstract these signals into a compact set of actionable principles (e.g., capability boundaries, safe/unsafe contexts, interface caveats, and common failure patterns) that guide future routing decisions. By continuously learning these routing principles as the agent pool evolves, the router becomes less prone to brittle dispatching caused by unfamiliarity with newly integrated agents, enabling end-to-end performance to scale more stably with additional agents. On the theory side, we cast orchestration under sequential augmentation as a contextual bandit and prove that our trust-region memory optimization yields a monotonic non-decreasing performance guarantee across onboarding rounds.
@@
  • Theoretical guarantee: under a contextual bandit formulation, we prove that our trust-region memory optimization yields a monotonic non-decreasing performance guarantee during scaling, providing a verifiable lower-bound safety property for open MAS expansion.
@@ 2 Related Works
LLM-based multi-agent systems (MAS) (Yang et al., 2025a; Chen et al., 2025; Yang et al., 2024) typically rely on a central router or coordinator to decompose tasks (Hong et al., 2024; Li et al., 2023; Wu et al., 2023), dispatch subtasks to specialized agents (Qian et al., 2024; Hu et al., 2025) (e.g., retrieval, tool use, code execution), and aggregate results (Wang et al., 2023, 2024), enabling tool-augmented and long-horizon workflows. In this setting, a large body of work studies how to improve planning, division of labor, and routing decisions under a static agent pool (Yue et al., 2025; Zhang et al., 2024; Liu et al., 2025b), i.e., the agent set and its capability profile are assumed fixed between training and deployment. Such approaches are effective when the action space is stable, but they usually do not provide a controlled online updating under continual agent augmentation mechanism that prevents regressions once the agent pool changes over time.
@@
Recent work explores self-evolving LLM-based multi-agent systems, where teams, roles, or interaction topologies are generated, optimized, or reorganized automatically (Yang et al., 2025a; Zhuge et al., 2024; Zhang et al., 2025a; Lu et al., 2025). These approaches demonstrate that MAS can improve via internal adaptation, but they often do not explicitly model the deployment reality of dynamic scale-up: continuously integrating new specialist or generalist agents into an existing system. Related work studies system-level expansion in LLM deployments by routing over growing pools of heterogeneous models (Wang et al., 2025; Zhang et al., 2025b). However, most prior work focuses on model-level routing among relatively stable LLMs. Agent routing in MAS is much more difficult: agents are far more heterogeneous (tools, memory, skills, roles), and often less reliable in practice (tool/API failures, brittle interfaces, retrieval noise), so routing must account for both capability matching and reliability rather than only cost-quality trade-offs.
@@
In LLM-based agent systems, a large body of work explores improving agent performance via diverse forms of learning. One common direction introduces explicit memory modules that store, organize, and update long-term experience (Zhang et al., 2026; Zhou et al., 2025; Zhang et al., 2025c; Ouyang et al., 2025), such as interaction trajectories, tool-use outcomes, and failure cases—so that agents can retrieve and reuse salient experience in subsequent decisions, enabling continual improvement and self-correction. Another line of research takes a policy-optimization perspective and leverages reinforcement learning and its variants to improve agent behaviors and alignment (Shao et al., 2024; Su et al., 2026; Dong et al., 2025); for example, GRPO has been used as an efficient policy-optimization algorithm for training LLMs, while recent training-free variants, such as Training-Free GRPO (Cai et al., 2025) shift optimization from parameter updates to contextual, experience-conditioned priors or policy adjustments at inference time, improving behaviors without updating model weights. More broadly, classic reinforcement learning ideas on conservative updates and stable improvement remain highly promising for agent learning (Schulman et al., 2017; Vieillard et al., 2020; Laroche et al., 2019). In this work, we model the MAS orchestration process as a contextual bandit and perform policy updates via memory learning during dynamic expansion, thereby reducing the risk of overall performance degradation or even performance collapse due to agent augmentation, while preserving existing capabilities as much as possible.
@@ 3.2.1 Expansion-Aware Task Synthesis
Our design draws on recent work on open-world agent trajectory synthesis (Guo et al., 2024; Fang et al., 2025), but adapts the philosophy from benchmark construction to generating a warm-up task distribution centered on new agents for safe Router updates.
@@ 3.2.2 Memory Update
The resulting evidence set (including both successful and failed routing cases) is aggregated to propose candidate memory edits. Concretely, we summarize the warm-up interactions into an experience buffer $\mathcal{B}_{k}=\{(x_{i},y_{i},r_{i})\}_{i=1}^{N_{k}}$, and distill auditable routing principles from $\mathcal{B}_{k}$ to form a set of candidate memory states (see full schema in Appendix D). Let $\tilde{\mathcal{U}}_{k}(m_{t})\subseteq\mathcal{M}$ denote the evidence-induced candidate set produced from $\mathcal{B}_{k}$. During candidate construction, we enforce a semantic trust region: newly distilled entries must be compatible with existing routing principles (e.g., via explicit applicability conditions, priorities, or exception clauses). Candidates with irreconcilable conflicts are rejected and thus do not enter $\tilde{\mathcal{U}}_{k}(m_{t})$. Note that $\tilde{\mathcal{U}}_{k}(m_{t})$ can be empty in degenerate cases where no consistent candidate can be distilled from the evidence.
@@ 4 Theoretical Analysis
In this section, we provide a monotonic non-decreasing performance guarantee across expansion stages for our memory update protocol under the contextual bandit formulation.
@@ 4.2 Monotonic Improvement Across Expansions
Under the non-interfering expansion assumption (Appendix), the conservative lift preserves performance:
@@
Let $\{m_{k}\}_{k\geq 0}$ be produced by (16). Then for all $k\geq 1$,
@@
See Appendix B.4. ∎
Table 1: Performance comparison on GAIA. MonoScale (using Qwen-3) shows strict monotonic improvement, eventually outperforming comparable strong baselines.
@@
MonoScale Scaling
Qwen-3-30B-A3B-Instruct(3 Agents) 49.06% 47.67% 26.92% 44.84%
5 Agents (w/ Memory) 52.83% 45.35% 42.31% 47.27% $\uparrow$+5.4%
7 Agents (w/ Memory) 60.38% 47.67% 30.77% 49.09% $\uparrow$+9.5%
@@ 5.2 Main Results
We first examine the scalability of our approach on the GAIA benchmark. As illustrated in Figure 2, MonoScale demonstrates a monotonic improvement. Table 1 quantifies this gain: as the agent pool expands from 3 to 10, the Qwen-3 based system steadily improves its overall accuracy from 44.84% to 55.15%, confirming that the router effectively integrates new capabilities without suffering from the cold-start confusion of expanding action spaces.
@@
In terms of absolute capability, MonoScale empowers a smaller open-weight model to rival proprietary giants. As shown in the lower section of Table 1, our 10-agent system (55.15%) outperforms the 3-agent baseline of DeepSeek-V3.2 (51.52%) and is competitive with GPT-5 (54.55%) and Gemini-2.5-Pro (54.55%) on GAIA. This advantage extends to expert-level reasoning tasks as well. Table 2 shows that on the challenging HLE benchmark, MonoScale improves accuracy from 11.60% to 19.90%, surpassing the static GPT-5 baseline (19.63%). These results demonstrate that MonoScale can enable a scaled MAS to outperform fixed-pool MAS configurations orchestrated by significantly larger SOTA models. Additional results and analysis could be found in Appendix C.1, and detailed case study could be found in Appendix E.
Table 2: Performance comparison on HLE. MonoScale (using Qwen-3) shows strict monotonic improvement with scaling, eventually surpassing the static GPT-5 baseline.
@@
  3 Agents, w/o Memory 11.60%
  5 Agents (w/ Memory) 15.60% $\uparrow$+34.5%
  7 Agents (w/ Memory) 18.70% $\uparrow$+61.2%
@@
Finally, we analyze robustness to unreliable workers. As illustrated in Figure 3, naive scale-up without memory updates can exhibit unstable, non-monotonic performance as the number of agents increases (even with a strong router such as Gemini-3-Pro). In contrast, MonoScale maintains a more stable scaling curve by proactively identifying failure modes during familiarization and encoding “negative constraints” into memory, which helps isolate malfunctioning agents during inference and improves reliability. Additional results and analysis could be found in Appendix C.2, and detailed case study could be found in Appendix E.
@@
6 Discussion
In the emerging Agentic Web (Yang et al., 2025b), a router may need to choose among millions of third-party agents uploaded by individuals or organizations. Even with basic quality control, agent cards can be incomplete, outdated, or misaligned with actual execution behavior, making zero-shot routing unreliable; this is especially problematic when tool interfaces are brittle, or tasks require multi-step workflows where a single misrouting can cascade into end-to-end failure. In this setting, simply upgrading to a stronger router model does not fundamentally address the lack of execution-grounded evidence. Instead, treating agent integration as a mandatory warm-up protocol, where we probe new agents with a small set of agent-conditioned tasks and distill both successes and failures into auditable, rollbackable routing constraints, appears essential for preventing post-expansion performance collapse.
Due to compute and experimental-scale constraints, we currently validate our approach only on relatively small agent pools of up to $10$ agents, where we observe strong effectiveness and stability. In realistic Agentic Web settings, routers face catalogs containing thousands to millions of agents, and routing is typically retrieval-driven: the system must first retrieve a small candidate set from a large directory before selecting and invoking an agent. How to continuously and in parallel onboard and expand the agent set at Web scale with a controllable cost budget—while warming up and calibrating routing to avoid degradation during expansion—remains an open challenge. A further question is whether retrieval-based routing should be more tightly integrated into our expansion strategy, forming a budgeted “retrieve–route–calibrate” loop that prioritizes calibration for newly surfaced or highly uncertain long-tail agents and feeds the resulting reliability and boundary information back into subsequent retrieval and routing decisions. Designing a more efficient, scalable, and stable expansion mechanism along these lines is a central focus of our future work.

### v2

@@ Abstract
In recent years, LLM-based multi-agent systems (MAS) have advanced rapidly, using a router to decompose tasks and delegate subtasks to specialized agents. A natural way to expand capability is to scale up the agent pool by continually integrating new functional agents or tool interfaces, but naive expansion can trigger performance collapse when the router cold-starts on newly added, heterogeneous, and unreliable agents. We propose MonoScale, an expansion-aware update framework that proactively generates a small set of agent-conditioned familiarization tasks, harvests evidence from both successful and failed interactions, and distills it into auditable natural-language memory to guide future routing. We formalize sequential augmentation as a contextual bandit and perform trust-region memory updates, yielding a monotonic non-decreasing performance guarantee across onboarding rounds under a non-interfering expansion assumption. Experiments on GAIA and Humanity’s Last Exam show stable gains as the agent pool grows, outperforming naive scale-up and strong-router fixed-pool baselines. Our code is available here.
@@ 1 Introduction
Multi-agent systems (MAS) built on large language models (LLMs) are emerging as a powerful paradigm for solving complex tasks (Chen et al., 2025; Yang et al., 2024; Guo et al., 2024; Liao et al., 2026). Unlike a single model that produces an answer end-to-end, an MAS typically includes a router that decomposes an input problem, dispatches subtasks to specialized agents (Hu et al., 2025; Yang et al., 2025b; Yue et al., 2025)(e.g., retrieval, code execution, planning, tool use, or domain experts), and then aggregates their outputs into a final response. This modular “divide–and–conquer” structure enables flexible collaboration and makes MAS particularly effective for multi-step workflows, tool-augmented reasoning, and long-horizon task execution.
@@
Existing approaches for improving multi-agent coordination broadly fall into two lines, yet neither fully addresses the central challenge of open-ended MAS scaling. The first line focuses on a static agent pool, assuming that the agent set and its capability profile remain stable between training and deployment (Yue et al., 2025; Zhang et al., 2024; Liu et al., 2025b); the research emphasis is then on optimizing routing policies, improving division of labor, or planning more efficiently within a fixed pool. While effective in static settings, these methods typically lack a controlled update mechanism that can prevent performance regressions when new agents are continually added, and they do not provide a unified formalization for online updating under continual agent augmentation in open-ended systems. In parallel, some recent work has started to examine dynamic expansion in LLM routing (Wang et al., 2025; Zhang et al., 2025c), where the system routes across an enlarging set of available experts over time. However, scaling an MAS is substantially harder: newly integrated agents are far more heterogeneous (tools, memory, skills, roles, and interfaces), and their quality can be highly uneven in practice (brittle tool calls, API failures, retrieval noise, or unstable behaviors). As a result, the router faces a sharper cold start problem, and a fixed onboarding suite to “trial-run” the new agent is often insufficient: (i) it cannot be adapted to each agent’s unique boundaries and failure modes; and (ii) it may under-cover the true deployment distribution, creating a risk of agents that appear usable during onboarding but still collapse after real deployment. In short, what dynamic MAS scaling truly requires is an adaptive (agent-conditioned) and conservative update protocol, so that end-to-end performance can improve stably as the agent pool grows.
@@
To bridge this gap, we propose an update framework for sequentially expanding open-ended MAS. Whenever a new agent is introduced, rather than waiting for real user traffic to surface failures, we proactively customize a small set of familiarization tasks tailored to the new agent and use them to collect controlled feedback. Our mechanism has two stages: (i) active familiarization via task customization, where tasks are synthesized to probe the new agent’s likely strengths, interface constraints, and failure modes, enabling the router to learn when to use / when not to use the agent; and (ii) router update via memory, where we update the router’s policy by maintaining an auditable text memory distilled from experience. Concretely, the memory captures not only patterns behind successful orchestration, but also lessons from failed worker calls and mis-routings; we abstract these signals into a compact set of actionable principles (e.g., capability boundaries, safe/unsafe contexts, interface caveats, and common failure patterns) that guide future routing decisions. By continuously learning these routing principles as the agent pool evolves, the router becomes less prone to brittle dispatching caused by unfamiliarity with newly integrated agents, enabling end-to-end performance to scale more stably with additional agents. On the theory side, we cast orchestration under sequential augmentation as a contextual bandit and prove that, under a non-interfering expansion assumption, our trust-region memory optimization yields a monotonic non-decreasing performance guarantee across onboarding rounds.
@@
  • Theoretical guarantee: under a contextual bandit formulation, we prove that, under a non-interfering expansion assumption, our trust-region memory optimization yields a monotonic non-decreasing performance guarantee during scaling, providing a verifiable conditional lower-bound for open MAS expansion.
@@ 2 Related Works
LLM-based multi-agent systems (MAS) (Yang et al., 2025b; Chen et al., 2025; Yang et al., 2024) typically rely on a central router or coordinator to decompose tasks (Hong et al., 2024; Li et al., 2023; Wu et al., 2023), dispatch subtasks to specialized agents (Qian et al., 2024; Hu et al., 2025) (e.g., retrieval, tool use, code execution), and aggregate results (Wang et al., 2023, 2024), enabling tool-augmented and long-horizon workflows. In this setting, a large body of work studies how to improve planning, division of labor, and routing decisions under a static agent pool (Yue et al., 2025; Zhang et al., 2024; Liu et al., 2025b), i.e., the agent set and its capability profile are assumed fixed between training and deployment. Such approaches are effective when the action space is stable, but they usually do not provide a controlled online updating under continual agent augmentation mechanism that prevents regressions once the agent pool changes over time.
@@
Recent work explores self-evolving LLM-based multi-agent systems, where teams, roles, or interaction topologies are generated, optimized, or reorganized automatically (Yang et al., 2025b; Zhuge et al., 2024; Zhang et al., 2025a; Lu et al., 2025). Such internal adaptation can improve a MAS, but unconstrained self-evolution may drift into emergent risks (Shao et al., 2026), so the evolutionary capability and reliability of agents must be explicitly evaluated rather than assumed (Fu et al., 2026; Yang et al., 2025a). Crucially, these approaches also do not address dynamic scale-up—continuously integrating new specialist or generalist agents into an existing system. A related line studies system-level expansion by routing over growing pools of heterogeneous models (Wang et al., 2025; Zhang et al., 2025c): UniRoute routes queries to previously unseen LLMs via learned model embeddings (Jitkrittum et al., 2025), and StageRoute performs online deployment and routing over a streaming model inventory under capacity and cost budgets (Li and Li, 2026). However, these address model-level routing, where the decision unit is a single standalone LLM. Agent routing in MAS is harder: agents are far more heterogeneous (tools, memory, skills, roles) and less reliable in practice (tool/API failures, brittle interfaces, retrieval noise), so routing must jointly weigh capability and reliability and produce a multi-step orchestration plan rather than a single-model choice. Closest to our setting, EvoRoute performs experience-driven self-routing within an agentic system (Zhang et al., 2026a), but targets the cost, latency, and accuracy trade-off rather than preventing cold-start collapse during agent-pool expansion. These methods are thus complementary to MonoScale rather than directly applicable baselines.
@@
A large body of work improves LLM-based agents via diverse forms of learning. One direction maintains explicit memory modules that distill long-term experience—trajectories, tool-use outcomes, and failure cases—for reuse in later decisions (Zhang et al., 2026c; Zhou et al., 2025; Zhang et al., 2025d; Ouyang et al., 2025; Wang et al., 2026). Another line applies reinforcement learning and its variants to improve agent behavior (Shao et al., 2024; Su et al., 2026; Dong et al., 2025; Liao et al., 2025; Zhang et al., 2025b); recent training-free variants such as Training-Free GRPO (Cai et al., 2025) move optimization from weight updates to inference-time, experience-conditioned policy adjustments. A further line distills modular, reusable skills from experience that can be retrieved or composed for new tasks (Zhang et al., 2026b; Xia et al., 2026; Lu et al., 2026). More broadly, classic ideas on conservative, stable policy updates remain relevant for agent learning (Schulman et al., 2017; Vieillard et al., 2020; Laroche et al., 2019). We model MAS orchestration as a contextual bandit and update the router via memory during expansion, mitigating performance collapse from agent augmentation while preserving existing capabilities.
@@ 3.2.1 Expansion-Aware Task Synthesis
Our design draws on recent work on open-world agent trajectory synthesis (Guo et al., 2024; Fang et al., 2025; Shi et al., 2025), but adapts the philosophy from benchmark construction to generating a warm-up task distribution centered on new agents for safe Router updates.
@@ 3.2.2 Memory Update
The resulting evidence set (including both successful and failed routing cases) is aggregated to propose candidate memory edits. Concretely, we summarize the warm-up interactions into an experience buffer $\mathcal{B}_{k}=\{(x_{i},y_{i},r_{i})\}_{i=1}^{N_{k}}$, and distill auditable routing principles from $\mathcal{B}_{k}$ to form a set of candidate memory states (see full schema in Appendix E). Let $\tilde{\mathcal{U}}_{k}(m_{t})\subseteq\mathcal{M}$ denote the evidence-induced candidate set produced from $\mathcal{B}_{k}$. During candidate construction, we enforce a semantic trust region: newly distilled entries must be compatible with existing routing principles (e.g., via explicit applicability conditions, priorities, or exception clauses). Candidates with irreconcilable conflicts are rejected and thus do not enter $\tilde{\mathcal{U}}_{k}(m_{t})$. Note that $\tilde{\mathcal{U}}_{k}(m_{t})$ can be empty in degenerate cases where no consistent candidate can be distilled from the evidence.
@@
Practical instantiation.
The constrained objective above serves as a theoretical lens for the stability analysis in Section 4: in deployment we do not continuously optimize the router, but instead perform a one-shot conservative selection over a finite candidate set. Concretely, with the router frozen, we synthesize agent-conditioned warm-up tasks, collect success and failure rollouts into $\mathcal{B}_{k}$, distill a finite set of candidate memories, and select a safe update that always retains the conservative fallback $m_{t}^{\uparrow}$. Algorithm 1 in Appendix B details the full procedure.
@@ 4 Theoretical Analysis
In this section, we provide a monotonic non-decreasing performance guarantee across expansion stages for our memory update protocol under the contextual bandit formulation and a non-interfering expansion assumption (Assumption C.1).
@@ 4.2 Monotonic Improvement Across Expansions
Under the non-interfering expansion assumption (Assumption C.1), the conservative lift preserves performance:
@@
Under Assumption C.1 (non-interfering expansion) and Assumption C.2 (feasible conservative fallback), let $\{m_{k}\}_{k\geq 0}$ be produced by (16). Then for all $k\geq 1$,
@@
See Appendix C.5. ∎
Remark 4.4 (Scope of the guarantee).
Theorem 4.3 is a conditional guarantee rather than a universal one: it holds when expansion is non-interfering, i.e., adding a new agent leaves the reward of any plan that does not invoke it unchanged. This assumption may fail under persistent environment drift—e.g., changing external API behavior or availability—where legacy-plan rewards shift for reasons orthogonal to agent expansion; we empirically examine its validity in Appendix C.2.
Table 1: Performance comparison on GAIA. For the Qwen-3-30B-A3B-Instruct backbone we report matched w/o Memory (naive scale-up) and w/ Memory (MonoScale) results at each agent count; green arrows mark the w/ Memory gain over w/o Memory at the same count. MonoScale improves consistently with scale and eventually outperforms comparable strong baselines.
@@
Naive Scale-up (Qwen-3-30B-A3B-Instruct, w/o Memory)
3 Agents 49.06% 47.67% 26.92% 44.84%
5 Agents 54.72% 43.02% 23.08% 43.64%
7 Agents 54.72% 40.70% 46.15% 46.06%
10 Agents 47.17% 44.19% 26.92% 42.42%
MonoScale (Qwen-3-30B-A3B-Instruct, w/ Memory)
3 Agents 49.06% 47.67% 26.92% 44.84%
5 Agents 52.83% 45.35% 42.31% 47.27% $\uparrow$+8.3%
7 Agents 60.38% 47.67% 30.77% 49.09% $\uparrow$+6.6%
@@ 5.2 Main Results
We first examine the scalability of our approach on the GAIA benchmark. As illustrated in Figure 2, MonoScale demonstrates a clear upward scaling trend, with only a small local dip at $N=6$ that we analyze in Appendix D.7. Table 1 quantifies this gain: as the agent pool expands from 3 to 10, the Qwen-3 based system steadily improves its overall accuracy from 44.84% to 55.15%, confirming that the router effectively integrates new capabilities without suffering from the cold-start confusion of expanding action spaces.
@@
In terms of absolute capability, MonoScale empowers a smaller open-weight model to rival proprietary giants. As shown in the lower section of Table 1, our 10-agent system (55.15%) outperforms the 3-agent baseline of DeepSeek-V3.2 (51.52%) and is competitive with GPT-5 (54.55%) and Gemini-2.5-Pro (54.55%) on GAIA. This advantage extends to expert-level reasoning tasks as well. Table 2 shows that on the challenging HLE benchmark, MonoScale improves accuracy from 11.70% to 19.88%, surpassing the static GPT-5 baseline (19.63%). These results demonstrate that MonoScale can enable a scaled MAS to outperform fixed-pool MAS configurations orchestrated by significantly larger SOTA models. Additional results and analysis could be found in Appendix D.1, an ablation study in Appendix D.2, a cross-router memory-transfer study in Appendix D.3, analyses of scaling beyond ten agents, onboarding cost, and multi-seed variance in Appendices D.4D.6, and detailed case studies in Appendix F.
Table 2: Performance comparison on HLE. MonoScale (using Qwen-3) shows consistent improvement with scaling, eventually surpassing the static GPT-5 baseline.
@@
  3 Agents, w/o Memory 11.70%
  5 Agents (w/ Memory) 15.59% $\uparrow$+33.2%
  7 Agents (w/ Memory) 18.71% $\uparrow$+59.9%
@@
Finally, we analyze robustness to unreliable workers. As illustrated in Figure 3, naive scale-up without memory updates can exhibit unstable, non-monotonic performance as the number of agents increases (even with a strong router such as Gemini-3-Pro). In contrast, MonoScale maintains a more stable scaling curve by proactively identifying failure modes during familiarization and encoding “negative constraints” into memory, which helps isolate malfunctioning agents during inference and improves reliability. Additional results and analysis could be found in Appendix D.8, and detailed case study could be found in Appendix F.
@@
6 Discussion and Limitations
In the emerging Agentic Web (Yang et al., 2025c), a router may need to choose among millions of third-party agents uploaded by individuals or organizations. Even with basic quality control, agent cards can be incomplete, outdated, or misaligned with actual execution behavior, making zero-shot routing unreliable; this is especially problematic when tool interfaces are brittle, or tasks require multi-step workflows where a single misrouting can cascade into end-to-end failure. In this setting, simply upgrading to a stronger router model does not fundamentally address the lack of execution-grounded evidence. Instead, treating agent integration as a mandatory warm-up protocol, where we probe new agents with a small set of agent-conditioned tasks and distill both successes and failures into auditable, rollbackable routing constraints, appears essential for preventing post-expansion performance collapse.
A practical consequence of this collapse risk shapes our experimental comparison. We benchmark MonoScale mainly against strong-router baselines at a fixed small ($3$-agent) pool rather than against their best naively scaled configuration, because of an asymmetry in predictability: under naive scale-up a strong router’s performance is non-monotonic, and which pool size is best can only be identified in hindsight, after exhaustively evaluating every scale. A practitioner onboarding new agents thus has no reliable way to know a priori at how many agents such a router will perform well—or where it will collapse—so its fixed small-pool configuration is the only stable, a-priori-knowable operating point, and hence the meaningful reference. MonoScale removes this uncertainty: its conservative updates yield a stable, non-decreasing scaling trend by construction, so its scaled performance is predictable and one can keep onboarding agents without gambling on where the peak lies.
Limitations.
We note several limitations. First, the monotonic-improvement guarantee is conditional: it holds under the non-interfering expansion assumption (Assumption C.1) and may weaken under persistent environment drift, such as shifting external-API behavior, as examined empirically in Appendix C.2. Second, MonoScale stores routing knowledge in a natural-language memory under a fixed token budget; at the agent-pool sizes we study we observe no saturation-induced collapse or catastrophic interference, but this is not established for much longer expansion horizons, where explicit memory management—retrieval, pruning, or consolidation—would likely be needed.
Due to compute and experimental-scale constraints, our main experiments validate the approach on relatively small agent pools of up to $10$ agents (with a preliminary $20$-agent probe in Appendix D.4), where we observe strong effectiveness and stability. In realistic Agentic Web settings, routers face catalogs containing thousands to millions of agents, and routing is typically retrieval-driven: the system must first retrieve a small candidate set from a large directory before selecting and invoking an agent. How to continuously and in parallel onboard and expand the agent set at Web scale with a controllable cost budget—while warming up and calibrating routing to avoid degradation during expansion—remains an open challenge. A further question is whether retrieval-based routing should be more tightly integrated into our expansion strategy, forming a budgeted “retrieve–route–calibrate” loop that prioritizes calibration for newly surfaced or highly uncertain long-tail agents and feeds the resulting reliability and boundary information back into subsequent retrieval and routing decisions. Designing a more efficient, scalable, and stable expansion mechanism along these lines is a central focus of our future work.